# FILP-3D: Enhancing 3D Few-shot Class-incremental Learning with Pre-trained Vision-Language Models

### v1

@@ 3.3 PTM is a Good 3D FSCIL Learner
Recent works [20, 47] have shown that PTMs substantially enhance the performance in incremental learning tasks. However, Chowdhury et al. [3]’s method, in the absence of shape-related prior knowledge, struggles with the challenge of catastrophic forgetting in 3D continual learning. Further complicating matters, the point features extracted by PointNet [18] are not in alignment with the prototypes generated through word2vec [15]. Thus, Chowdhury et al. [3]’s work finds it challenging to simultaneously retain features associated with base classes (representing old knowledge) and integrate features of novel classes (representing new knowledge).
@@
In response to the aforementioned challenges, our initial approach was to embed shape-related prior knowledge via a 3D PTM. However, the substantial disparity in quality and volume between 3D data and its text/2D counterparts renders current 3D PTMs less effective in generalizing to downstream tasks. Motivated by this limitation, we turn to an alternative method. Drawing inspiration from recent studies [43, 11], we leverage CLIP [21] to indirectly infuse shape-related prior knowledge through projection. We christen this new framework as SimpleCIL-3D.
@@
Specifically, we project the point cloud data into multi-view depth maps $\mathbf{D}^{b}_{1:N}$. A pre-trained ViT [8] in CLIP is then deployed to extract depth features $\mathbf{F}^{2D}\in\mathbb{R}^{N\times C}$. Here, $N$ is the number of views, and $C$ is the embedding dimension of ViT. To allow incremental tasks, a learnable merger is attached, formulating global depth features $\mathbf{f}^{d}\in\mathbb{R}^{C}$ as follows,
$\mathbf{f}^{d}=f^{1}_{d}({\rm ReLU}(f^{2}_{d}({\rm concat}(\mathbf{F}_{1:N}^{2D}))))$ (1)
where $f^{1}_{d}$, $f^{2}_{d}$ are two learnable MLPs.
@@
The efficacy of setting classifier weights to average embeddings (referred to as prototypes) for CIL tasks has been well-established by [25]. However, while Chowdhury et al. [3]’s method employs word2vec to construct prototypes for new classes, it falls short in encapsulating the average semantics of 3D point clouds. In our approach, the depth features, denoted as $\mathbf{f}^{d}$, are pre-aligned with text embeddings of CLIP. This enables SimpleCIL-3D to progressively produce prototypes using a CLIP text encoder. For each class labeled as $k$ with its respective name $t_{k}$, we introduce a template text prompt: “an image or projection or sketch of a $t_{k}$”. This is mapped to the CLIP prototype symbolized by $\mathbf{F}^{t}_{k}\in\mathbb{R}^{C}$. Owing to the inherent association of our method between image and textual representations, there is no compulsion to realign the two modalities during incremental phases, unlike strategies such as in Chowdhury et al. [3]’s work. As a result, we can directly utilize the cosine similarity between $\mathbf{f}^{d}$ and $\mathbf{F}^{t}_{k}$ to represent the logit for class $k$. The ultimate probability prediction, represented by $\mathbf{p}$, is formulated as follows:
$l_{k}=\cos(\mathbf{f}^{d},\mathbf{F}^{t}_{k}),\quad\mathbf{p}={\rm softmax}([l_{1},\dots,l_{K}]).$ (2)
@@ 3.4 Ensure Performance of PTM in 3D FSCIL
Large-scale V-L PTMs undeniably serve as rich repositories of prior knowledge. However, their application to 3D FSCIL tasks reveals unique challenges. First, point cloud data, which consists of discrete points in 3D space, primarily represents geometric features of objects. It inherently misses richer visual nuances, such as color and texture. This deficiency is especially significant given CLIP’s visual encoder, which inherently seeks and incorporates these visual details into its semantic representation. Consequently, this could jeopardize the precise classification of depth maps. Second, point clouds derived from real-world scans often suffer from noise, which distorts the extraction of depth-centric features. This problem is accentuated in multi-view projection techniques, where a significant fraction of the points might be occluded, thereby increasing the potential noise. Such disturbances can upset the balance in few-shot incremental learning, leading to overfitting. Addressing these challenges is imperative when tailoring PTMs for 3D FSCIL tasks. In the ensuing sections, we aim to enhance SimpleCIL-3D with elucidate solutions tailored to these specific hurdles, to obtain FILP-3D.
@@ 3.4.1 Redundant Feature Eliminator
We propose our design to mitigate the adverse effect of superfluous features here. Our empirical observations indicate that redundant features typically exhibit minimal inter-class distinction, leading to constrained variance. By contrast, semantic features pertinent for classification distinctly express a pronounced bias for each class, resulting in a more pronounced variance.  Chowdhury et al. [3], Zhu et al. [48] also yield empirical observations akin to ours. Drawing upon this insight, we can discern between redundant and semantic features based on their variance. Subsequently, we employ Dimensionality Reduction (DR) techniques to preserve the semantic essence while condensing the extraneous features.
@@
Figure 3: The current feature space is generated by three principal components. Each dimension’s one-hot vector represents a principal component. The first two dimensions contain semantic information (high variance in the x and y axes), while the third dimension serves as a redundant component (low variance in the z-axis). By transforming feature vector (green) into the feature space mentioned above, we can notice that projection can eliminate redundant information, while normalization will improperly stretch the semantic information.
@@
Following the above discussions, we can obtain the following formulation for $\widetilde{l}_{k}$, i.e., the logit after eliminating superfluous features. For the redundant feature eliminated logit of the $k$-th class (the detail discussion is in Sec. 1 of Suppl.), we have:
$\widetilde{l}_{k}=\frac{\widetilde{\mathbf{f}}^{d}\widetilde{\mathbf{F}}^{t^{\mathrm{T}}}_{k}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$ (3)
where $\widetilde{\mathbf{f}}^{d}=\mathbf{V}^{\mathrm{T}}\mathbf{f}^{d}$ and $\widetilde{\mathbf{F}}^{t}_{k}=\mathbf{V}^{\mathrm{T}}\mathbf{F}^{t}_{k}$. Here, $\mathbf{V}=[\mathbf{v}_{i}]_{M}$ consists of the principal components extracted from the base task ($M$ is the number of principal components)
Note that, different from calculating the cosine similarity of $\widetilde{\mathbf{f}}^{d}$ and $\widetilde{\mathbf{F}}^{t}_{k}$, the denominator in Eq. 3 is the norm of the original features $\mathbf{f}^{d}$ and $\mathbf{F}^{t}_{k}$. Principal features are directly dot multiplied after normalizing with the original features, and such a procedure is named as the renormalized cosine similarity (RCS). We can qualitatively compare the above two similarity formulations: If the mode lengths of $\widetilde{\mathbf{f}}^{d}$ and $\widetilde{\mathbf{F}}^{t}_{k}$ are relatively small, it indicates that only a small portion of the features can be accurately extracted. Performing normalization under such circumstances will magnify the originally small features, resulting in incorrect classification, as shown in Figure 3. Conversely, the dot method has no such improper stretching operations.
@@ 3.4.2 Spatial Noise Compensator
SimpleCIL-3D unifies pre-aligned image and text encoders through comprehensive data training. This ensures a substantial reservoir of prior knowledge, minimizing the potential for information loss during any subsequent alignment processes. Given these attributes, it stands as a markedly superior choice when juxtaposed against partially-aligned 3D pre-trained models. Nonetheless, the projection-based methodology inherent to SimpleCIL-3D exhibits heightened sensitivity to both noise and viewpoint selection. When noise obscures the clarity of depth maps, rendering them ineffective in delineating object contours, there is a marked downturn in classification efficacy. As a countermeasure, we propose the incorporation of a graph-based 3D model. This addition is devised to augment the compromised information, bolstering our model’s resilience against noise disturbances.
The 3D module yields features that are invariant to transformations, thereby enhancing their robustness against noise. On the other hand, the multi-view model, fortified by PTM, extracts semantically richer and more encompassing information. To harness the strengths of both, we amalgamate the features derived from the 3D module and the multi-view model. The global feature is computed as:
$\mathbf{f}^{p}=f^{2}_{p}({\rm ReLU}(f^{1}_{p}(\mathbf{f}^{3D})))$ (4)
$\mathbf{f}^{g}=\frac{1}{2}({\rm max}(\mathbf{f}^{d},\mathbf{f}^{p})+{\rm avg}(\mathbf{f}^{d},\mathbf{f}^{p}))$ (5)
where $\mathbf{f}^{3D}$ is the feature extracted by 3D encoder from raw point cloud data $\mathbf{x}_{i}^{b}$, and $\mathbf{f}^{g}$ represents new global features used to replace $\mathbf{f}^{d}$ for further tasks. $f^{1}_{p}$, $f^{2}_{p}$ are learnable MLPs, so that 3D features $\mathbf{f}^{p}$ can be aligned to multi-view features $\mathbf{f}^{d}$.
Using this approach, any information inadequately captured due to viewpoint selection is supplemented by the 3D channel using the ${\rm max}$ operation. Concurrently, overly dominant features are moderated through the ${\rm average}$ operation for balance.
@@ 3.5 Training Objective
The parameters of the 3D encoder and the multi-view render & encoder are fixed during both base and incremental training stage. Only Merger (Eq. 1) for depth feature fusion and Adapter (Eq. 4) for 3D feature alignment are trained. The classification loss $\mathcal{L}^{b}_{cls}$ for $b$-th task can be calculated as follows:
@@
where $\mathcal{L}_{ce}$ is cross-entropy loss [4], $\mathbf{p}^{b}_{i}$ donates the predicted probability of the $i$-th sample of the $b$-th task.
Inspired by  [38, 26], we also utilize contrastive learning [1] and data augmentation in the training of 3D FSCIL models for continual learning. Through the push and pull dynamics of contrastive learning, we optimize the extracted features, bringing them closer to the appropriate prototype while distancing them from erroneous ones. This strategy notably reduces ambiguity throughout the continual learning phase. Concurrently, data augmentation not only amplifies the efficacy of contrastive learning but also acts as a safeguard against overfitting in few-shot scenarios.
@@
where $N_{Aug}$ is the number of augmentations, and $\mathbf{f}^{g}_{i,j}$ is the global feature encoded after replacing $x^{b}_{i}$ with $f_{Aug}(x^{b}_{i})$.
The overall training loss $\mathcal{L}^{b}$ for $b$-th task can be calculated as follows:
$\mathcal{L}^{b}=\mathcal{L}^{b}_{cls}+\alpha\mathcal{L}^{b}_{cont}$ (8)

### v2

@@ 4.3 PTM is a Good 3D FSCIL Learner
Recent works [11, 16] have shown that PTMs substantially enhance performance in incremental learning tasks. However, Chowdhury et al. [8]’s method lacks shape-related prior knowledge and naturally aligned features/text encoders, thus struggling to retain old knowledge while learning new knowledge.
@@
In response to the aforementioned challenges, we attempt to introduce PTM to incorporate prior knowledge. Our initial approach was to embed shape-related prior knowledge via a 3D PTM [23]. However, the substantial disparity in quality and volume between 3D data [27, 13] and its text/2D counterparts [12] renders current 3D PTMs less effective in generalizing to downstream tasks, like zero/few-shot classification [17]. Motivated by this limitation, we turn to an alternative method. Drawing inspiration from recent studies [15, 17], we leverage CLIP [12] to indirectly infuse shape-related prior knowledge through projection. We christen this new framework as our baseline.
@@
Specifically, we project the point cloud data into multi-view depth maps. A pre-trained ViT [1] in CLIP is then deployed to extract depth features $\mathbf{F}^{d}\in\mathbb{R}^{N\times C}$. Here, $N$ is the number of views, and $C$ is the embedding dimension of ViT. To allow incremental tasks, a learnable merger is attached, formulating global 2D features $\mathbf{f}^{2D}\in\mathbb{R}^{C}$ as follows,
$\mathbf{f}^{2D}={\rm E}(\Phi({\rm F}^{d}))$ (1)
where $\rm E$ is a two-layer perceptron with ReLU activation function and $\Phi$ denotes the concat operation. This simple adapter can effectively act as a Merger to merge multi-view features into a global feature [17].
@@
The efficacy of setting classifier weights to average embeddings (referred to as prototypes) for CIL tasks has been well-established by [28]. However, while Chowdhury et al. [8]’s method employs word2vec to construct prototypes for new classes, it falls short in encapsulating the average semantics of 3D point clouds. In our approach, the 2D features are pre-aligned with text embeddings of CLIP. This enables our baseline to progressively produce prototypes using a CLIP text encoder. For each class labeled as $k$ with its respective name $t_{k}$, we introduce a template text prompt: “an image or projection or sketch of a $t_{k}$”. This is mapped to the CLIP prototype symbolized by $\mathbf{F}^{t}_{k}\in\mathbb{R}^{C}$. Owing to the inherent association of our method between image and textual representations, there is no compulsion to realign the two modalities during incremental phases, unlike strategies such as in Chowdhury et al. [8]’s work. As a result, we can directly utilize the cosine similarity between $\mathbf{f}^{2D}$ and $\mathbf{F}^{t}_{k}$ to represent the logit for class $k$. The ultimate probability prediction, represented by $\mathbf{p}$, is formulated as follows:
$l_{k}=\cos(\mathbf{f}^{2D},\mathbf{F}^{t}_{k}),\quad\mathbf{p}={\rm softmax}([l_{1},\dots,l_{K}]).$ (2)
@@
Despite these improvements, the incompatibility between 3D data representation and the 2D feature encoder impede further performance enhancements, primarily manifested as two aspects.
@@ 4.4 Ensure Performance of PTM in 3D FSCIL
Large-scale V-L PTMs undeniably serve as rich repositories of prior knowledge. However, their application to 3D FSCIL tasks reveals unique challenge: mismatch between mismatch between 3D representation and V-L PTM’s encoder. This mismatch is primarily manifested in two aspects: feature space misalignment and significant noise. In the ensuing sections, we aim to enhance the baseline with elucidating solutions tailored to these specific hurdles, to obtain FILP-3D. Specifically, we propose the Redundant Feature Eliminator (RFE) to address the feature space misalignment issue and the Spatial Noise Compensato (SNC) to tackle the problem of significant noise.
@@ 4.4.1 Redundant Feature Eliminator
We propose our design to mitigate the adverse effect of feature space misalignment here. Our empirical observations indicate that redundant features typically exhibit minimal inter-class distinction, leading to constrained variance. By contrast, semantic features pertinent for classification distinctly express a pronounced bias for each class, resulting in a more pronounced variance.  Chowdhury et al. [8], Zhu et al. [9] also yield empirical observations akin to ours. Drawing upon this insight, we can discern between redundant and semantic features based on their variance. Subsequently, we employ Dimensionality Reduction (DR) techniques to preserve the semantic essence while condensing the extraneous features.
@@
Figure 3: The current feature space is generated by three principal components. Each dimension’s one-hot vector represents a principal component. The first two dimensions contain semantic information, while the third dimension serves as a redundant component. By transforming feature vector (green) into the feature space mentioned above, we can notice that projection can eliminate redundant information, while normalization will improperly stretch the semantic information, leading to misclassification.
@@
Following the above discussions, we can obtain the formulation for calculating $\widetilde{l}_{k}$, $i.e.$, the logit after eliminating superfluous features. The detailed discussion is as follows.
$\mathbf{V}$ is the principal components extracted from the base task, $\mathbf{v}_{i}$ is the $i$-th normalized principal component. $\widetilde{\mathbf{f}}^{2D}$ and $\widetilde{\mathbf{F}}^{t}_{k}$ are the projection of $\mathbf{f}^{2D}$ and $\mathbf{F}^{t}_{k}$ onto the $\mathbf{V}$ vector space, formally expressed as the following equation:
$\mathbf{f}^{2D}=\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{2D}_{i}\mathbf{v}_{i}+\mathbf{R}^{2D}\ \mathbf{F}^{t}_{k}=\sum_{i=1}^{M}\widetilde{\mathbf{F}}^{t}_{k,i}\mathbf{v}_{i}+\mathbf{R}^{t}_{k}$ (3)
where $M$ is the number of principal components, $\widetilde{\mathbf{f}}^{2D}_{i}$ and $\mathbf{F}^{t}_{k,i}$ are the $i$-th dimension of $\widetilde{\mathbf{f}}^{2D}$ and $\widetilde{\mathbf{F}}^{t}_{k}$, respectively, and $\mathbf{R}^{2D}$ and $\mathbf{R}^{t}_{k}$ are the components of $\mathbf{f}^{2D}$ and $\mathbf{F}^{t}_{k}$ which are not in the $\mathbf{V}$ vector space. $\mathbf{R}^{2D}$ and $\mathbf{R}^{t}_{k}$ are low-variance components (due to the nature of principal component analysis) and thus are redundant components according to previous discussions.
Each $\mathbf{v}_{i}$ is an orthogonal vector and orthogonal to vectors outside the $\mathbf{V}$ vector space, e.g., $\mathbf{R}^{2D}_{k}$ and $\mathbf{R}^{t}_{k}$. Therefore $\forall i,j,k,\ i\neq j$ The following equation holds:
$\mathbf{v}_{i}\mathbf{v}^{T}_{j}=0,\ \mathbf{v}_{i}\mathbf{v}^{T}_{i}=1,\ \mathbf{R}^{2D}_{k}\mathbf{v}_{i}^{T}=0,\ \mathbf{v}_{i}\mathbf{R}^{t^{T}}_{k}=0$ (4)
Then, we can derive the following equation:
$l_{k}=\frac{\mathbf{f}^{2D}\mathbf{F}^{t^{\mathrm{T}}}_{k}}{\|\mathbf{f}^{2D}\|\|\mathbf{F}^{t}_{k}\|}$ $=\frac{(\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{2D}_{i}\mathbf{v}_{i}+\mathbf{R}^{2D})(\sum_{i=1}^{M}\widetilde{\mathbf{F}}^{t}_{k,i}\mathbf{v}_{i}+\mathbf{R}^{t}_{k})^{T}}{\|\mathbf{f}^{2D}\|\|\mathbf{F}^{t}_{k}\|}$
$=\frac{\widetilde{\mathbf{f}}^{2D}\widetilde{\mathbf{F}}^{t^{T}}_{k}+\mathbf{R}^{2D}\mathbf{R}^{t^{T}}_{k}}{\|\mathbf{f}^{2D}\|\|\mathbf{F}^{t}_{k}\|}$
Now we can observe that there are two terms in the numerator of this equation. $\widetilde{\mathbf{f}}^{2D}\widetilde{\mathbf{F}}^{t^{T}}_{k}$ represents semantic information and should be retained, and $\mathbf{R}^{2D}\mathbf{R}^{t^{T}}_{k}$ is the multiplication of redundant feature vectors that should be compressed to keep only semantic information. Thus, the logit of the class $k$ can be calculated in a modified way as follows to eliminate redundant features:
$\widetilde{l}_{k}=\frac{\widetilde{\mathbf{f}}^{2D}\widetilde{\mathbf{F}}^{t^{T}}_{k}}{\|\mathbf{f}^{2D}\|\|\mathbf{F}^{t}_{k}\|}.$ (5)
Note that, different from calculating the cosine similarity of $\widetilde{\mathbf{f}}^{2D}$ and $\widetilde{\mathbf{F}}^{t}_{k}$, the denominator in Eq. 5 is the norm of the original features $\mathbf{f}^{2D}$ and $\mathbf{F}^{t}_{k}$. Principal features are directly dot multiplied after normalizing with the original features, and such a procedure is named as the renormalized cosine similarity (RCS). As shown in Figure 3. Both the direction and length of the projected vectors influence the classification result in RCS. However, the regularization operation in direct cosine similarity disregards the projected vector’s length, which leads to incorrect classification. For example: The class of the input sample is airplane and the wrong class is table. The x and y axes denote the presence or absence of wings and the shape of the backbone. The z-axis denote the surface texture. The airplane has wings, an oval backbone, and metal surface, so the yellow and green arrows are high in the x and y axes and low in the z-axis. The table does no have a wing and has a rectangle backbone and wooden surface, so the purple arrow is low in the x and y axes and high in the z-axis. After projection, the yellow and green arrows are high along both the x and y axes, while the purple arrow is low along both axes, allowing the sample to be correctly classified as an airplane. However, after normalization, the purple arrow becomes high along both the x and y axes as well, causing the sample to be misclassified as a table.
Details of Pre-Processing.
Figure 4: Overview of pre-processing
In the pre-processing stage, we render all training samples of the base task $\{\mathbf{x}^{1}_{1},...,\mathbf{x}^{1}_{M}\}$ to generate their depth maps $\{\rm\mathbf{D}^{1}_{1},...,\mathbf{D}^{1}_{M}\}$ and then embed them as depth features $\rm\mathbf{F}^{d}\in\mathbb{R}^{(|\mathcal{D}^{1}|N)\times C}$ by the ViT in CLIP, where $|\mathcal{D}^{1}|$ is the number of training sample in task 1, $N$ is the number of viewpoint and $C$ is the embedding dimension. We then use the depth features ${\rm\mathbf{F}^{2D}}$ to calculate principal components $\mathbf{V}\in\mathbb{R}^{M\times C}$ through SVD, as is shown in Figure 4, where $M$ is the number of principal components.
@@ 4.4.2 Spatial Noise Compensator
Baseline unifies pre-aligned image and text encoders through comprehensive data training. This ensures a substantial reservoir of prior knowledge, minimizing the potential for information loss during any subsequent alignment processes. Given these attributes, it stands as a markedly superior choice when juxtaposed against partially aligned 3D pre-trained models. Nonetheless, the projection-based methodology inherent to the baseline exhibits heightened sensitivity to both noise and viewpoint selection. When noise obscures the clarity of depth maps, rendering them ineffective in delineating object contours, there is a marked downturn in classification efficacy. The supervised classification performance of other projection-based methods like PointCLIP [15] (72.8% Acc.) and CLIP2Point [17] (74.5%) compared to DGCNN [21] (78.1%) and PointMAE [24] (86.4%) on ScanObjectNN also verify ours viewpoint. As a countermeasure, we propose the incorporation of a graph-based 3D model. This addition is devised to augment the compromised information, bolstering our model’s resilience against noise disturbances.
The 3D module yields features that are invariant to transformations, thereby enhancing their robustness against noise. On the other hand, the multi-view model, fortified by PTM, extracts semantically richer and more encompassing information. To harness the strengths of both, we amalgamate the features derived from the 3D module and the multi-view model.
Following ULIP [3], FILP-3D adapts cross-modal contrastive learning to achieve alignment between 3D and CLIP text/image. This alignment operation maps the features of 3D, image, and text to the same feature space, enabling a more rational fusion of 2D and 3D features and effective classification. The align loss $\mathcal{L}_{(M1,M2)}$ between modal $M1,M2$ and the final align loss $\mathcal{L}_{align}$ can be calculated as follows:
$\mathcal{L}_{(M1,M2)}=\sum_{(i,j)}-\frac{1}{2}log\frac{\exp\left(\frac{\mathbf{h}_{i}^{M_{1}}\mathbf{h}_{j}^{M_{2}}}{\tau}\right)}{\sum_{k}\exp\left(\frac{\mathbf{h}_{i}^{M_{1}}\mathbf{h}_{k}^{M_{2}}}{\tau}\right)}-\frac{1}{2}log\frac{\exp\left(\frac{\mathbf{h}_{i}^{M_{1}}\mathbf{h}_{j}^{M_{2}}}{\tau}\right)}{\sum_{k}\exp\left(\frac{\mathbf{h}_{k}^{M_{1}}\mathbf{h}_{j}^{M_{2}}}{\tau}\right)}$ (6)
$\mathcal{L}_{align}=\alpha\mathcal{L}_{(P,S)}+\beta\mathcal{L}_{(I,S)}$ (7)
where $\mathbf{h}_{i}^{M}$ donates the $i$-th feature in modal $M$, $\tau$ donates the temperature parameter and the $P,I,S$ donate the set of point cloud modal $\mathbf{f}^{3D}$, the set of image modal $\mathbf{f}^{2D}$ and text modal $\mathbf{F}^{t}$ in this session respectively. $\mathbf{f}^{3D}$ can be calculated as follows:
$\mathbf{f}^{3D}={\rm E}({\rm\mathbf{f}^{p}})$ (8)
where $\rm E$ is a two-layer perceptron with ReLU activation function and is named as the Adapter.
After the alignment, we utilize the mixed pooling [29] to fuse 2D features and 3D features. According to [29], mixed pooling is superior to single max/average pooling in terms of feature fusion. Using this approach, any information inadequately captured due to viewpoint selection is supplemented by the 3D channel using the ${\rm max}$ operation. Concurrently, overly dominant features are moderated through the ${\rm average}$ operation for balance. The new global feature can be calculated as follows:
$\mathbf{f}^{g}=\frac{1}{2}({\rm max}(\mathbf{f}^{2D},\mathbf{f}^{3D})+{\rm average}(\mathbf{f}^{2D},\mathbf{f}^{3D}))$ (9)
where $\mathbf{f}^{g}$ represents global features used to replace $\mathbf{f}^{2D}$ for further tasks.
@@ 4.5 Training Objective
The parameters of the 3D encoder and the multi-view render & encoder are fixed during both base and incremental training stages. Only Merger (Eq. 1) for depth feature fusion and Adapter (Eq. 8) for 3D feature alignment are trained. The classification loss $\mathcal{L}^{b}_{cls}$ for $b$-th task can be calculated as follows:
@@
where $\mathcal{L}_{ce}$ is cross-entropy loss [30], $\mathbf{p}^{b}_{i}$ denotes the predicted probability of the $i$-th sample of the $b$-th task.
Inspired by  [31], we also utilize contrastive learning [32] and data augmentation in the training of 3D FSCIL models for continual learning. Through the push and pull dynamics of contrastive learning, we optimize the extracted features, bringing them closer to the appropriate prototype while distancing them from erroneous ones. This strategy notably reduces ambiguity throughout the continual learning phase. Concurrently, data augmentation not only amplifies the efficacy of contrastive learning but also acts as a safeguard against overfitting in few-shot scenarios.
@@
where $N_{Aug}$ is the number of augmentation iterations, and $\mathbf{f}^{g}_{i,j}$ is the global feature encoded after replacing $x^{b}_{i}$ with $f_{Aug}(x^{b}_{i})$.
The overall training loss $\mathcal{L}^{b}$ for $b$-th task can be calculated as the sum of align loss Eq 7, classification loss10 and contrastive learning loss 11 follows:
$\mathcal{L}^{b}=\mathcal{L}^{b}_{align}+\mathcal{L}^{b}_{cls}+\mathcal{L}^{b}_{cont}$ (12)